# BARI-TH/25-772

### v1

@@ 3 Observables
The couplings (21)-(34) show that in the ABCD model flavour changing neutral current transitions occur at tree-level through $Z^{\prime}$ exchanges, and that lepton flavour violating processes are possible. For this reason, we study the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition, both for $\ell_{1}=\ell_{2}$ and $\ell_{1}\neq\ell_{2}$, focusing on the exclusive modes $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$, $B\to K^{*}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$ for which a wealth of experimental information is available and some tensions are spotted. Moreover, we describe the leptonic LFV modes $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, and the $Z^{\prime}$ contribution to $(g-2)_{\mu}$.
@@ 3.1 Effective Hamiltonian for $b\to s\ell_{1}^{-}\ell_{2}^{+}$
[…]
$G_{F}$ is the Fermi constant, $V_{ij}$ are elements of the CKM matrix, and doubly Cabibbo suppressed terms proportional to $V_{ub}V_{us}^{*}$ are neglected. [^7] The complete basis of operators including (pseudo-)scalar and tensor operators can be found in [8]. The effective Hamiltonian includes the current-current operators $O_{1,2}$
@@
$O_{5}$ $=$ $({\bar{s}}\gamma^{\mu}P_{L}\,b)\sum_{q}({\bar{q}}\gamma^{\mu}P_{R}\,q)\,\,,\quad O_{6}=({\bar{s}}_{\alpha}\gamma^{\mu}P_{L}\,b_{\beta})\sum_{q}({\bar{q}}_{\beta}\gamma^{\mu}P_{R}\,q_{\alpha})\,\,,$ (41)
with $P_{R,L}=\displaystyle\frac{1\pm\gamma_{5}}{2}$, $\alpha$,$\beta$ colour indices, and the sum in (40)-(41) running over the flavours $q=u,d,s,c,b$. The other operators in (38) are the magnetic penguin operators
@@
In SM, at the leading order in $\alpha_{s}$ the only operators to be considered are $O_{7}$ and $O_{9,10}$ with $\ell_{1}=\ell_{2}$. At this order, $O_{7}$ is the only operator contributing to $b\to s\gamma$. The renormalization group evolution to the scale $\mu_{b}\simeq{\cal O}(m_{b})$ also involves the magnetic gluon penguin operator $O_{8}$ and the operators $O_{1,\dots 6}$. Their mixing into $O_{7}$ generates large logarithms producing a strong enhancement of the rates. The anomalous dimension matrix governing the mixing is regularization scheme dependent. One can get rid of such a dependence defining an effective coefficient $C_{7}^{(0)\rm eff}(\mu_{b})$ which includes contributions of $O_{1,\dots 6}$ [27]. Non factorizable contributions to the matrix elements of $O_{1}$ and $O_{2}$ must also be taken into account. Their effects produce a further shift in the coefficient of $O_{7}$: $C_{7}^{\rm eff}=1.33\,C_{7}^{(0)\rm eff}$ [28]. In this way, $O_{7}$ turns out to give the dominant contribution to $b\to s\gamma$, with the SM Wilson coefficients known at NNLO in QCD [29, 2]. $O_{7}$ contributes to $b\to s\ell^{+}_{1}\ell^{-}_{2}$ for $\ell_{1}=\ell_{2}$, due to the photon conversion into the lepton pair.
In SM, the contribution to $b\to s\ell^{+}\ell^{-}$ comes from the operators $O_{9}$ and $O_{10}$ and also from other operators in (38), in particular $O_{1}$ and $O_{2}$ due to charm re-scattering. This contribution overwhelms the contributions of the other operators in the region of $q^{2}$, the dilepton invariant mass, corresponding to the narrow charmonium resonances $J/\psi$, $\psi(2S)$, …The contribution of charm re-scattering in the other intervals of $q^{2}$, the role of which has been pointed out long ago [30], has theoretical uncertainties difficult to evaluate. A possibility to account for it is to replace $C_{9}$ with $C_{9}^{\rm eff}=C_{9}+Y(q^{2})$ [31, 32, 33, 34, 35], introducing a function $Y(q^{2})$ which comprises the terms described in [36]. This could obscure genuine new short-distance contributions to $C_{9}$. Charm re-scattering contributions can be described through non-local form factors determined, e.g., by dispersion relations employing information from the nonleptonic modes $B\to K^{*}(J/\psi,\psi(2S))$ [37, 38, 39, 36], with an uncertainty difficult to estimate [40]. In the comparison with measurements it is customary to exclude the $q^{2}$ region corresponding to the resonances $J/\psi$, $\psi(2S)$, which are cut in the experimental analyses to isolate the short-distance part of the $q^{2}$ spectrum [41]. Since charm re-scattering does not affect the LFV processes with $\ell_{1}\neq\ell_{2}$, of prime interest in our study, for the $\ell^{+}\ell^{-}$ mode we also cut the resonance region and omit to include charm rescattering effects in the remaining $q^{2}$ interval.
@@ 3.1.1 $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$
For the decay $B_{s}(p)\to\ell_{1}^{-}(k_{1})\ell_{2}^{+}(k_{2})$, with $p,k_{1,2}$ the particle four-momenta, a single hadronic quantity is required, the decay constant $f_{B_{s}}$ defined by
@@
¯${\bar{\cal B}}(B_{s}\to\ell^{-}_{1}\ell_{2}^{+})$ $=$ $\frac{\tau_{B_{s}}}{(1-y_{s})}\frac{G_{F}^{2}\alpha^{2}|\lambda_{ts}^{*}|^{2}f_{B_{s}}^{2}}{64\pi^{3}m_{B_{s}}^{3}}\lambda^{1/2}(m_{B_{s}}^{2},\,m_{\ell_{1}}^{2},\,m_{\ell_{2}}^{2})$ (49)
@@
with $\lambda_{ts}^{*}=V_{tb}V_{ts}^{*}$, $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ the Källen function, $\tau_{B_{s}}$ the average $B_{s}$ lifetime. The notation ${\bar{\cal B}}$ indicates that the effects of the $B_{s}-\bar{B}_{s}$ oscillations are taken into account through the factor $1/(1-y_{s})$ [42], with $y_{s}=0.061\pm 0.009$ [43]. In SM with $\ell_{1}=\ell_{2}$ Eq. (49) shows that only the operator $O_{10}$ contributes to this decay.
@@ 3.1.2 $\bar{B}^{0}\to\bar{K}^{*0}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$
The function $I(q^{2},\theta_{\ell},\theta_{K^{*}},\phi)$ can be written in terms of angular coefficient functions $I_{i}^{(a)}(q^{2})$ using the decomposition [8]
@@
and on the angular $P$-observables [44]
$P_{1}(q^{2})$ $=$ $\displaystyle\frac{2}{F_{T}(q^{2})}S_{3}(q^{2})\,\,,\quad P_{2}(q^{2})=\displaystyle\frac{1}{2F_{T}(q^{2})}S_{6s}(q^{2})\,\,,\quad P_{3}(q^{2})=-\displaystyle\frac{1}{F_{T}(q^{2})}S_{9}(q^{2})\,\,,\quad$
@@
The forward-backward asymmetry is recognised to be sensitive to NP, in particular because in SM it changes sign in the physical $q^{2}$ range [45, 46, 47, 35, 48], with the zero $q_{0}^{2}$ determined by
@@
A model independent prediction for $q_{0}^{2}$ is obtained in the large energy limit of $K^{*}$, due to spin symmetry relations among the form factors [49]
@@
where $E$ is the $K^{*}$ energy in the $B$ rest frame. [^8] The relations (59) are broken by hard gluon corrections to the weak vertex and by hard spectator interactions [50]. Since $m_{K^{*}}^{2}\ll m_{B}^{2}$, Eqs. (58),(59) give
@@
is sensitive to the virtual photon polarization [51, 52], which in SM is mainly left-handed, the right-handed polarization being suppressed by $m_{s}/m_{b}$. Deviations are expected in scenarios where the photon can also be right-handed, an effect encoded in the operator $O_{7}^{\prime}$, which is not generated in the scenario A of the ABCD model.
@@ 3.2 LFV lepton decays
At present, the LFV modes $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ are under intense theoretical and experimental scrutiny. Without neutrino oscillations such processes are absent in SM. In a minimal extension allowing for neutrino masses and mixing, they can occur at rates much smaller than the present experimental upper bounds discussed below.
@@
    In a version of the SM extended with the see-saw mechanism to describe massive nutrinos this process can occur though a penguin diagram with internal $W$ and oscillating neutrinos. The photon is emitted by $W$. If there is a $Z^{\prime}$ with LFV couplings, other penguin diagrams contribute, with $Z^{\prime}$ and a charged lepton in the loop. The branching fraction is [53]
@@
    $[O_{VXY}]^{ij}_{kl}=(\bar{\psi}_{i}\gamma^{\mu}P_{X}\psi_{j})(\bar{\psi}_{k}\gamma_{\mu}P_{Y}\psi_{l})\,\,,\qquad[C_{VXY}]^{ij}_{kl}=\frac{\Delta^{ij}_{X}\Delta^{kl}_{Y}}{M^{2}_{Z^{\prime}}}\,\,,$ (66)
    with $X,Y=L,R$. The branching fraction of $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ is given by [1]:

### v2

@@ 3 Observables
The couplings (21)-(34) show that in the ABCD model flavour changing neutral current transitions occur at tree-level through $Z^{\prime}$ exchanges, and that lepton flavour violating processes are possible. For this reason, we study the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition, both for $\ell_{1}=\ell_{2}$ and $\ell_{1}\neq\ell_{2}$, focusing on the exclusive modes $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$, $B\to K^{*}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$ for which a wealth of experimental information is available and some tensions are spotted. Moreover, we describe the leptonic LFV modes $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}e^{+}e^{-}$ and the $\mu^{-}\to e^{-}$ conversion in nuclei. Finally, we consider the $Z^{\prime}$ contribution to $(g-2)_{\mu}$.
@@ 3.1 Effective Hamiltonian for $b\to s\ell_{1}^{-}\ell_{2}^{+}$
$H^{\rm eff}=-\,4\,\frac{G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}\,$ $\Big\{C_{1}O_{1}+C_{2}O_{2}+\sum_{i=3,..,6}C_{i}O_{i}+\sum_{i=7,..,10}\left[C_{i}O_{i}+C_{i}^{\prime}O_{i}^{\prime}\right]\Big\}\,\,.$ (38)
$G_{F}$ is the Fermi constant, $V_{ij}$ are elements of the CKM matrix, and doubly Cabibbo suppressed terms proportional to $V_{ub}V_{us}^{*}$ are neglected. [^6] The complete basis of operators including (pseudo-)scalar and tensor operators can be found in [Altmannshofer:2008dz]. The effective Hamiltonian includes the current-current operators $O_{1,2}$
@@
$O_{5}$ $=$ $({\bar{s}}\gamma^{\mu}P_{L}\,b)\sum_{q}({\bar{q}}\gamma_{\mu}P_{R}\,q)\,\,,\quad O_{6}=({\bar{s}}_{\alpha}\gamma^{\mu}P_{L}\,b_{\beta})\sum_{q}({\bar{q}}_{\beta}\gamma_{\mu}P_{R}\,q_{\alpha})\,\,$, (41)
with $P_{R,L}=\displaystyle\frac{1\pm\gamma_{5}}{2}$, $\alpha$$\beta$ colour indices, and the sum in (40)-(41) running over the flavours $q=u,d,s,c,b$. The other operators in (38) are the magnetic penguin operators
@@
In SM, at the leading order in $\alpha_{s}$ the only operators to be considered are $O_{7}$ and $O_{9,10}$ with $\ell_{1}=\ell_{2}$. At this order, $O_{7}$ is the only operator contributing to $b\to s\gamma$. The renormalization group evolution to the scale $\mu_{b}\simeq{\cal O}(m_{b})$ also involves the magnetic gluon penguin operator $O_{8}$ and the operators $O_{1,\dots 6}$. Their mixing into $O_{7}$ generates large logarithms producing a strong enhancement of the rates. The anomalous dimension matrix governing the mixing is regularization scheme dependent. One can get rid of such a dependence defining an effective coefficient $C_{7}^{(0)\rm eff}(\mu_{b})$ which includes contributions of $O_{1,\dots 6}$ [Buras:1993xp]. Non factorizable contributions to the matrix elements of $O_{1}$ and $O_{2}$ must also be taken into account. Their effects produce a further shift in the coefficient of $O_{7}$: $C_{7}^{\rm eff}=1.33\,C_{7}^{(0)\rm eff}$ [Beneke:2001at]. In this way, $O_{7}$ turns out to give the dominant contribution to $b\to s\gamma$, with the SM Wilson coefficients known at NNLO in QCD [Misiak:2018cec, Buras:2020xsm]. $O_{7}$ contributes to $b\to s\ell^{+}_{1}\ell^{-}_{2}$ for $\ell_{1}=\ell_{2}$, due to the photon conversion into the lepton pair.
In SM, the contribution to $b\to s\ell^{+}\ell^{-}$ comes from the operators $O_{9}$ and $O_{10}$ and also from other operators in (38), in particular $O_{1}$ and $O_{2}$ due to charm re-scattering. This contribution overwhelms the contributions of the other operators in the region of $q^{2}$, the dilepton invariant mass, corresponding to the narrow charmonium resonances $J/\psi$, $\psi(2S)$, …The contribution of charm re-scattering in the other intervals of $q^{2}$, the role of which has been pointed out long ago [Colangelo:1989gi], has theoretical uncertainties difficult to evaluate. A possibility to account for it is to replace $C_{9}$ with $C_{9}^{\rm eff}=C_{9}+Y(q^{2})$ [Lim:1988yu, Deshpande:1988bd, ODonnell:1991cdx, Paver:1991tn, Colangelo:1995jv], introducing a function $Y(q^{2})$ which comprises the terms described in [Bordone:2024hui]. This could obscure genuine new short-distance contributions to $C_{9}$. Charm re-scattering contributions can be described through non-local form factors determined, e.g., by dispersion relations employing information from the nonleptonic modes $B\to K^{*}(J/\psi,\psi(2S))$ [Khodjamirian:2010vf, Khodjamirian:2012rm, Bobeth:2017vxj, Bordone:2024hui], with an uncertainty difficult to estimate [Ciuchini:2022wbq]. In the comparison with measurements it is customary to exclude the $q^{2}$ region corresponding to the resonances $J/\psi$, $\psi(2S)$, which are cut in the experimental analyses to isolate the short-distance part of the $q^{2}$ spectrum [LHCb:2024onj]. Since charm re-scattering does not affect the LFV processes with $\ell_{1}\neq\ell_{2}$, of prime interest in our study, for the $\ell^{+}\ell^{-}$ mode we also cut the resonance region and omit to include charm rescattering effects in the remaining $q^{2}$ interval.
@@ 3.1.1 $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$
For the decay $B_{s}(p)\to\ell_{1}^{-}(k_{1})\ell_{2}^{+}(k_{2})$, with $p$ and $k_{1,2}$ the particle four-momenta, a single hadronic quantity is required, the decay constant $f_{B_{s}}$ defined by
@@
${\bar{\cal B}}(B_{s}\to\ell^{-}_{1}\ell_{2}^{+})$ $=$ $\frac{\tau_{B_{s}}}{(1-y_{s})}\frac{G_{F}^{2}\alpha^{2}|\lambda_{ts}^{*}|^{2}f_{B_{s}}^{2}}{64\pi^{3}m_{B_{s}}^{3}}\lambda^{1/2}(m_{B_{s}}^{2},\,m_{\ell_{1}}^{2},\,m_{\ell_{2}}^{2})$ (49)
@@
with $\lambda_{ts}^{*}=V_{tb}V_{ts}^{*}$, $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ the Källen function, $\tau_{B_{s}}$ the average $B_{s}$ lifetime. The notation $\bar{\cal B}$} indicates that the effects of the $B_{s}-\bar{B}_{s}$ oscillations are taken into account through the factor $1/(1-y_{s})$ [DeBruyn:2012wj], with $y_{s}=0.061\pm 0.009$ [LHCb:2014iah]. In SM with $\ell_{1}=\ell_{2}$ Eq. (49) shows that only the operator $O_{10}$ contributes to this decay.
@@ 3.1.2 $\bar{B}^{0}\to\bar{K}^{*0}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$
The function $I(q^{2},\theta_{\ell},\theta_{K^{*}},\phi)$ can be written in terms of angular coefficient functions $I_{i}^{(a)}(q^{2})$ using the decomposition [Altmannshofer:2008dz]
@@
and on the angular $P$-observables [Matias:2012xw]
$P_{1}(q^{2})$ $=$ $\displaystyle\frac{2}{F_{T}(q^{2})}S_{3}(q^{2})\,\,,\quad P_{2}(q^{2})=\displaystyle\frac{1}{2F_{T}(q^{2})}S_{6s}(q^{2})\,\,,\quad P_{3}(q^{2})=-\displaystyle\frac{1}{F_{T}(q^{2})}S_{9}(q^{2})\,\,$$\quad$
@@
The forward-backward asymmetry is recognised to be sensitive to NP, in particular because in SM it changes sign in the physical $q^{2}$ range [Ali:1991is, Ali:1994bf, Liu:1994cfa, Colangelo:1995jv, Ali:1999mm], with the zero $q_{0}^{2}$ determined by
@@
A model independent prediction for $q_{0}^{2}$ is obtained in the large energy limit of $K^{*}$, due to spin symmetry relations among the form factors [Charles:1998dr]
@@
where $E$ is the $K^{*}$ energy in the $B$ rest frame. [^7] The relations (59) are broken by hard gluon corrections to the weak vertex and by hard spectator interactions [Beneke:2000wa]. Since $m_{K^{*}}^{2}\ll m_{B}^{2}$, Eqs. (58),(59) give
@@
is sensitive to the virtual photon polarization [Kruger:2005ep, Becirevic:2011bp], which in SM is mainly left-handed, the right-handed polarization being suppressed by $m_{s}/m_{b}$. Deviations are expected in scenarios where the photon can also be right-handed, an effect encoded in the operator $O_{7}^{\prime}$, which is not generated in the scenario A of the ABCD model.
@@ 3.2 LFV lepton decays
In this section we examine the contribution of $Z^{\prime}$ to LFV leptonic modes. We begin with the processes $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ and $\mu^{-}\to e^{-}e^{+}e^{-}$. In absence of neutrino oscillations such processes do not occur in SM. In a minimal extension allowing for neutrino masses and mixing, they can occur at rates much smaller than the present experimental upper bounds discussed below. We then turn to the $\mu^{-}\to e^{-}$ conversion in nuclei.
@@
    In a version of the SM extended with the see-saw mechanism to describe massive neutrinos this process can occur through a penguin diagram with internal $W$ and oscillating neutrinos. The photon is emitted by $W$. If there is a $Z^{\prime}$ with LFV couplings, other penguin diagrams contribute, with $Z^{\prime}$ and a charged lepton in the loop. The branching fraction is [Lindner:2016bgg]
@@
    $O_{VXY}]^{ij}_{kl}=(\bar{\psi}_{i}\gamma^{\mu}P_{X}\psi_{j})(\bar{\psi}_{k}\gamma_{\mu}P_{Y}\psi_{l})\,\,,\qquad[C_{VXY}]^{ij}_{kl}=\frac{\Delta^{ij}_{X}\Delta^{kl}_{Y}}{M^{2}_{Z^{\prime}}}\,\,$, (66)
    with $X,Y=L,R$. The branching fraction of $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ is given by [Aebischer:2019blw]:
@@
  • $\mu^{-}-e^{-}$ conversion in nuclei
    The $\mu^{-}-e^{-}$ conversion in a nucleus of atomic number $Z$ and mass number $A$ is the process
    $\mu^{-}+(A,Z)\rightarrow e^{-}+(A,Z)\,.$ (68)
    The transition rate $\Gamma(\mu\rightarrow e)$ in the ABCD model is obtained considering the dominant tree-level $Z^{\prime}$ contributions inducing four-fermion operators, and assuming that the conversion leaves the nucleus in its ground state [Hisano:1995cp, Kitano:2002mt]:
    $\Gamma(\mu^{-}\rightarrow e^{-})$ $=$ $\frac{\alpha^{3}}{16\pi^{2}}\frac{Z_{eff}^{4}}{Z}|F(q)|^{2}~\frac{m_{\mu}^{5}}{M^{4}_{Z^{\prime}}}\left[|\Delta_{L}^{\mu e}(Z^{\prime})|^{2}+|\Delta_{R}^{\mu e}(Z^{\prime})|^{2}\right]$ (69)
    $\times\left|(2Z+N)\Delta_{V}^{uu}(Z^{\prime})+(Z+2N)\Delta_{V}^{dd}(Z^{\prime})\right|^{2}.$
    $\Delta^{ij}_{V}(Z^{\prime})$ are in (65). N=AZN=A-Z is the number of neutrons in the nucleus, $Z_{eff}$ is an effective parameter and $F(q^{2})$ the nuclear form factor. The observable to investigate is the branching fraction defined as
    $\mathcal{B}(\mu^{-}\to e^{-})=\frac{\Gamma(\mu^{-}\to e^{-})}{\Gamma_{\rm capt}}\,\,$, (70)
    which involves the muon capture rate $\Gamma_{\rm capt}$ in the nucleus. The upper bound for $\mathcal{B}(\mu\to e)$ in $^{48}_{22}{\rm{Ti}}$ established by the SINDRUM II experiment at PSI [Kaulard:1998rb] is
    $\mathcal{B}(\mu^{-}\to e^{-})\leq 4.3\times 10^{-12}\quad({\rm at}\,\,\,90\%\,\,{\rm C.L.})\,\,.$ (71)
    For $^{48}_{22}{\rm{Ti}}$ the values of the relevant parameters are $Z_{eff}=17.6$ and $F(q^{2}\simeq-m_{\mu}^{2})\simeq 0.54$ [Bernabeu:1993ta], with the muon capture rate $\Gamma_{\rm capt}\simeq(2.590\pm 0.012)\times 10^{6}$ s-1 [Suzuki:1987jf]. This is the case considered in our analysis. An upper bound has been established by the SINDRUM II Collaboration also for the gold target: $\mathcal{B}(\mu^{-}\,{\rm Au}\to e^{-}\,{\rm Au})\leq 7\times 10^{-13}$[SINDRUMII:2006dvw]. Its analysis requires a different set of nuclear input parameters, that can be inferred from [Kitano:2002mt, Suzuki:1987jf]. The results of our study do not change if the experimental bound and the parameters for the gold target are used.